OpenStudy (anonymous):

Show that the function y(t)=3e ^-t ^2/2 is a solution to the differential equation y' = -ty

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@satellite73 @jim_thompson5910

OpenStudy (anonymous):

Is the equation\[t(t)=3e^{ \dfrac{-t^{-t}}{2} } \]

OpenStudy (anonymous):

Woah that went wrong \[t(t)=3e^{ \dfrac{-t^{2}}{2} }\]

OpenStudy (anonymous):

yes that right

OpenStudy (anonymous):

y(t)*

OpenStudy (anonymous):

@tom982 do u know how todo this?

OpenStudy (anonymous):

Ah okay, y(t) makes way more sense. It's asking to show that y' = -ty so we need to find the derivative of y(t) and show that it's equal to -t*y(t). Can you tell me the derivative of y(t)?

OpenStudy (anonymous):

is it -t2/2 3e^1/2?

OpenStudy (anonymous):

-3e ^(-t^2/2 ) t

OpenStudy (loser66):

done

OpenStudy (anonymous):

@Loser66 ??

OpenStudy (anonymous):

Yep. The derivative is \[y'(t)=-3te^{\dfrac{-t^2}{2}}\]What do we get when we multiply\[y(t)=3e^{ \dfrac{-t^{2}}{2} }\] by -t?

OpenStudy (anonymous):

3e t^3/2?

OpenStudy (anonymous):

No it doesn't go into the power \[-t\times y(t)=-t\times3e^{ \dfrac{-t^{2}}{2} }=-3te^{ \dfrac{-t^{2}}{2}}=y'(t)\]As required.

OpenStudy (anonymous):

okay so that's the direvitive

OpenStudy (anonymous):

Yep, we've just shown that -ty(t)=y'(t) which is what the question wanted.

OpenStudy (anonymous):

the question asks for y' =-ty why do u have or why did u said (t)??

OpenStudy (anonymous):

is that the same thing?

OpenStudy (anonymous):

?? @tom982

OpenStudy (anonymous):

same y can write like this \[y(t)\] because y is function of t

OpenStudy (anonymous):

okat thanks I got it :) ty everyone for helping